Calculate the pH of the following solutions.

(a) 0.73 M NaOH

(b) 8.7 ✕ 10-10 M NaOH

(c) 2.5 M NaOH

Please show calculation and explain

I have calculated (a) is 13.86 and (c) is 14.4. Just need the answer for (b)
Answer #1

B) 8.7*10^-10 M NaOH

When concentration of soecies involved is so low then we consider contribution from water as well.

For eg. Water is composed of H+ and OH- ions

and [H+][OH-] = 10^-14

So [H+] =[OH-] = 10^-7

Total [OH-] in the solution is 10^-7 + 8.7*10^-10

[OH-] = 10^-10(10^3 + 8.7)

[OH-] = 10^-10(1000+8.7)

[OH-] = 1008.7*10^-10

POH = -log( 1008.7*10^-10)

POH = 6.99

So ph = 14- poh

Ph = 7.00 is the answer